### v1

@@ 6.1 Proof of results in Section 2
(6.2) $=Y_{t}^{x}+{\mathcal{}E}^{g}_{t}\bigg[-\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds+\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=:Y_{t}^{x}+\operatorname{II}_{t},$
where the equality holds by the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [12, Lemma 2.1].
Since it holds that $g(s,Z_{s}^{x})\leq|g(s,Z_{s}^{x})|\leq\kappa|Z_{s}^{x}|$ for all $s\in[t,\tau]$ (see Definition 2.1 (ii), (iii)), by the monotonicity of ${\mathcal{}E}_{t}^{g}[\cdot]$ (see [12, Proposition 2.2 (iii)]),
(6.3) $\operatorname{II}_{t}\leq{\mathcal{}E}_{t}^{g}\bigg[\kappa\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}|Z_{s}^{x}|ds+\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=:\operatorname{III}_{t}.$
We note that ${\mathcal{}E}^{g}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\to\mathbb{R}$ given in Definition 2.1 is an $\mathcal{}F$}-expectation [^5] A nonlinear expectation ${\mathcal{}E}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\to\mathbb{R}$ is called $\mathcal{}F$}-expectation if for each $\xi\in L^{2}({\mathcal{}F}_{T};\mathbb{R})$ and $t\in[0,T]$ there exists a random variable $\eta\in L^{2}({\mathcal{}F}_{t};\mathbb{R})$ such that ${\mathcal{}E}[\xi{\bf 1}_{A}]={\mathcal{}E}[\eta{\bf 1}_{A}]$ for all $A\in{\mathcal{}F}_{t}$. Moreover, given $\mu>0$, we say that an $\mathcal{}F$}-expectation $\mathcal{}E$} is dominated by ${\mathcal{}E}^{\mu}$ if for all $\xi,\eta\in L^{2}({\mathcal{}F}_{T};\mathbb{R})$ ${\mathcal{}E}(\xi+\eta)-{\mathcal{}E}(\xi)\leq{\mathcal{}E}^{\mu}[\eta];$ see [12, Definitions 3.2 and 4.1].. Moreover, by [12, Remark 4.1] it is dominated by a $g$-expectation ${\mathcal{}E}^{\kappa}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\to\mathbb{R}$ which is defined by setting that $g(\omega,t,z):=\kappa|z|$ for all $(\omega,t,z)\in\Omega\times[0,T]\times\mathbb{R}^{d}$, where the constant $\kappa>0$ appears in Definition 2.1 (ii).
Hence, an application of [12, Lemma 4.4] ensures that
(6.4) $\operatorname{III}_{t}\leq{\mathcal{}E}_{t}^{\kappa}\bigg[\kappa\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}|Z_{s}^{x}|ds+\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=0,$
where the equality holds because $(e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x})_{s\in[t,T]}$ is $\mathbb{F}$-predictable and satisfies $\mathbb{E}[\int_{t}^{T}|e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}|^{2}ds]<\infty$ (noting that $Z^{x}\in\mathbb{L}^{2}(\mathbb{R}^{d})$ and $\beta_{t}\geq 0$ for all $t\in[0,T]$; see Definition 2.6 and Assumption 2 (ii)), hence the integrand given in (6.4) is ${\mathcal{}E}^{\kappa}$-martingale and the corresponding $g$-expectation equals zero; see [12, Lemma 5.5].
Combining (6.2), (6.3) and (6.4), we obtain that ${\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}]\leq Y_{t}^{x}$ $\mathbb{P}$-a.s.. Since $\tau\in{\mathcal{}T}_{t}$ is chosen some arbitrary, we have $V_{t}^{x}=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}]\leq Y_{t}^{x}.$
Step 2. We now claim that $Y_{t}^{x}\leq V_{t}^{x}$. Let $\tau_{t}^{*,x}\in{\mathcal{}T}_{t}$ be defined as in (2.5). Since $\int_{0}^{\tau_{t}^{*,x}}(Y_{s-}^{x}-R(X_{s-}^{x}))dK_{s}^{x}=0$ $\mathbb{P}$-a.s. (see Definition 2.6 (iv)) and $Y_{s-}^{x}>R(X_{s-}^{x})$ for all $s\in(0,\tau_{t}^{*,x})$ (by definition of $\tau_{t}^{*,x}$), it holds that
@@
By putting $\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds-\int_{t}^{\tau_{t}^{*,x}}(e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x})^{\top}dB_{s}$ into the left-hand side of (6.6) and taking the conditional $g$-expectation ${\mathcal{}E}_{t}^{g}[\cdot]$, $\mathbb{P}$-a.s.,
(6.7) $\begin{aligned} \operatorname{III}_{t}^{x}&:={\mathcal{}E}_{t}^{g}\bigg[\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}r(X_{s}^{x})ds+e^{-\int_{t}^{\tau_{t}^{*,x}}\beta_{u}du}Y_{\tau^{*}}^{x}\bigg]\\ &\;=Y_{t}^{x}+{\mathcal{}E}_{t}^{g}\bigg[-\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds+\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]\\ &\;=:Y_{t}^{x}+\operatorname{IV}_{t}^{x},\end{aligned}$
where we have used the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [12, Lemma 2.1].
Since $Y_{\tau_{t}^{*,x}}^{x}\leq R(X_{\tau_{t}^{*,x}}^{x})$ on $\{\tau_{t}^{*,x}<T\}$; $Y_{\tau_{t}^{*,x}}^{x}=R(X_{\tau_{t}^{*,x}}^{x})$ on $\{\tau_{t}^{*,x}=T\}$, we have
(6.8) $\begin{aligned} \operatorname{III}_{t}^{x}&\leq{\mathcal{}E}_{t}^{g}\bigg[\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}r(X_{s}^{x})ds+e^{-\int_{t}^{\tau_{t}^{*,x}}\beta_{u}du}R(X_{\tau_{t}^{*,x}}^{x})\bigg]={\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}],\end{aligned}$
where $\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}\in L^{2}({\mathcal{}F}_{\tau^{*}};\mathbb{R})$ is given in (2.2) (under the setting $\tau=\tau_{t}^{*,x}$) and the last inequality follows from the positiveness of $(\beta_{u})_{u\in[0,T]}$.
Let ${\mathcal{}E}^{-\kappa}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\to\mathbb{R}$ be a $g$-expectation defined by setting $g(\omega,t,z):=-\kappa|z|$ for all $(\omega,t,z)\in\Omega\times[0,T]\times\mathbb{R}^{d}$. Then since it holds that $g(s,Z_{s}^{x})\geq-|g(s,Z_{s}^{x})|\geq-\kappa|Z_{s}^{x}|$ for all $s\in[t,\tau_{t}^{*,x}]$ (see Definition 2.1 (ii), (iii)),
(6.9) $\begin{aligned} \operatorname{IV}_{t}^{x}&\geq{\mathcal{}E}_{t}^{g}\bigg[-\kappa\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}|Z_{s}^{x}|ds+\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]\\ &\geq{\mathcal{}E}_{t}^{-\kappa}\bigg[-\kappa\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}|Z_{s}^{x}|ds+\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=0,\end{aligned}$
where the first inequality follows from the monotonicity of ${\mathcal{}E}_{t}^{g}[\cdot]$ (see [12, Proposition 2.2 (iii)]), the second inequality follows from [12, Lemma 4.4], and the last equality follows from the same arguments presented for the equality given in (6.4).
Combining (6.7)–(6.9), we obtain that $Y_{t}^{x}\leq{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}]$, $\mathbb{P}$-a.s.. As $\tau_{t}^{*,x}=\inf\{s\geq t\,|\,Y_{s}^{x}\leq R(X_{s}^{x})\}\wedge T\in{\mathcal{}T}_{t}$, we have $Y_{t}^{x}\leq V_{t}^{x}=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}],$ $\mathbb{P}$-a.s., as claimed. Therefore, $\tau_{t}^{*,x}$ given in (2.5) is optimal to (2.2). This completes the proof.
@@
Since $\alpha$ is uniformly bounded (noting that it has values only in $\{0,1\}$), one can deduce that the parameters of the BSDE (6.11) satisfies all the conditions given in [49, Section 3]. Hence, there exists a unique solution $(\widetilde{Y}_{t}^{x;N,\alpha},\widetilde{Z}_{t}^{x;N,\alpha})_{t\in[0,T]}\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ to the controlled BSDE (6.11).
@@
where we have used the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [12, Lemma 2.1].
Moreover, by using the same arguments presented for the ${\mathcal{}E}^{g}$-supermartingale property in (6.2)–(6.4) and the ${\mathcal{}E}^{g}$-submartingale property in (6.7) and (6.9) (see the proof of Proposition 2.8) we can deduce that the conditional $g$-expectation appearing in the right-hand side of the above equals zero (i.e., the integrand therein is an ${\mathcal{}E}^{g}$-martingale). Hence the claim holds.
Step 2. It suffices to show that for every $t\in[0,T]$ $\mathbb{P}$-a.s., $Y_{t}^{x;N}=\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{Y}_{t}^{x;N,\alpha}.$ Indeed, it follows from Step 1 that for every $\alpha\in{\mathcal{}A}$ the parameters of the BSDE (6.11) satisfies the conditions given in [49, Section 3]. Furthermore, the parameters of the BSDE (2.7) also satisfies the conditions (see Remark 2.9 (i)).
@@
(6.12) $F^{x;N}_{t}\big(Y_{t}^{x;N},Z_{t}^{x;N}\big)\geq\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{F}_{t}^{x;N,\alpha}(Y_{t}^{x;N},Z_{t}^{x;N}).$
@@
Therefore, an application of [21, Proposition 3.1] ensures the claim to hold.
Step 3. Lastly, it follows from [21, Corollary 3.3] that the process $\alpha^{*,x;N}\in{\mathcal{}A}$ is optimal for the problem given in Step 2., i.e., for all $t\in[0,T]$ $\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{Y}_{t}^{x;N,\alpha}=\widetilde{Y}_{t}^{x;N,\alpha^{*,x;N}}.$ This completes the proof.
@@ 6.2 Proof of results in Section 3
Let $N\in\mathbb{N}$ and $\lambda>0$ be given. We prove (i) by showing that the parameters of the BSDE (3.8) satisfy all the conditions given in [49, Section 3] to ensure its existence and uniqueness to hold.
@@
is (strictly) decreasing and $N\lambda^{-1}$-Lipschitz continuous, we are able to see that for every $\omega\in\Omega$, $t\in[0,T]$, and $y,\hat{y}\in\mathbb{R}$
@@
From (6.13) and (6.15) and the definition of $\overline{F}^{x;N,\lambda}$ given in (3.7), it follows that the desired priori estimate of $\overline{F}^{x;N,\lambda}$ holds. Hence an application of [49, Theorem 3.1] ensures the existence and uniqueness of the solution of (3.8), as claimed.
@@
(6.16) $\overline{F}_{t}^{x;N,\lambda}(\omega,y,z)=F_{t}^{x}(\omega,y,z)+\max_{a\in[0,1]}\bigg\{N(R(X_{t}^{x}(\omega))-y)a-\lambda{\mathcal{}H}(a)\bigg\},$
@@
Therefore, an application of [21, Proposition 3.1] ensures the claim to hold.
Moreover, a direct application of [21, Corollary 3.3] ensures that $\pi^{*,x;N,\lambda}$ is optimal for $\overline{V}^{x;N,\lambda}$ given in (3.2). This completes the proof.
@@
Moreover, as the terminal conditions of (3.8) and (2.7) are coincide, it follows from the comparison principle of BSDEs (see, e.g., [21, Theorem 2.2]) that (3.10) holds.
@@
Since the parameters of the BSDEs (3.8) and (2.7) satisfy the conditions given in [21, Section 5] (with exponent $2$) for all $N\in\mathbb{N}$ and $\lambda>0$, we are able to apply [21, Proposition 5.1] to have the following a priori estimates: [^6] In [21, Section 5], the filtration (denoted by $({\mathcal{}F}_{t})$ therein) is set to be right-continuous and complete (and hence not necessarily the Brownian filtration, as in our case). Nevertheless, we can still apply the stability result given in [21, Proposition 5.1], since the martingales $M^{i}$, $i=1,2$, appearing therein are orthogonal to the Brownian motion. Consequently, the arguments remain valid when the general filtration is replaced with the Brownian one. for every $N\in\mathbb{N}$ and $\lambda>0$
@@
Proof 6.5 (Proof of Corollary 3.6).
Set for every $N\in\mathbb{N}$ and $\lambda>0$, $D_{t}^{x;N}:=Y_{t}^{x;N}-R(X_{t}^{x})$ and $\overline{D}_{t}^{x;N,\lambda}:=\overline{Y}_{t}^{x;N,\lambda}-R(X_{t}^{x})$, $t\in[0,T]$, where $Y^{x;N}$ and $\overline{Y}^{x;N,\lambda}$ denote the first components of the unique solution to the BSDEs (2.7) and (3.8), respectively (see also Remark 2.9 and Theorem 3.4 (i)).

### v2

@@ 6.1 Proof of results in Section 2
(6.2) $=Y_{t}^{x}+{\mathcal{}E}^{g}_{t}\bigg[-\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds+\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]$
$=Y_{t}^{x}+{\mathcal{}E}^{g}_{t}\bigg[-\int_{t}^{\tau}g(s,e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x})ds+\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=:Y_{t}^{x}+\operatorname{II}_{t},$
where the first equality holds by the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [coquet2002filtration, Lemma 2.1]. The second equality holds since $0<e^{-\int_{t}^{s}\beta_{u}du}\leq 1$ (as $\beta_{t}\geq 0$ for all $t\in[0,T]$; Assumption 2 (ii)), and for every $\omega\in\Omega$ and $t\in[0,T]$, it holds that $g(\omega,t,cz)=cg(\omega,t,z)$ for all $z\in\mathbb{R}^{d}$ and $c\in(0,1)$; see Definition 2.1 (iii).
Moreover, since $Z^{x}\in\mathbb{L}^{2}(\mathbb{R}^{d})$, $(\widetilde{Z}_{s}^{x})_{s\in[t,T]}:=(e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x})_{s\in[t,T]}$ is $\mathbb{F}$-predictable and satisfies $\mathbb{E}[\int_{t}^{T}|\widetilde{Z}_{s}^{x}|^{2}ds]<\infty$. Therefore, the integrand of $\operatorname{II}_{t}$ is ${\mathcal{}E}^{g}$-martingale and its conditional $g$-expectation $\operatorname{II}_{t}$ equals zero; see [coquet2002filtration, Lemma 5.5].
Combining this with (6.2), we obtain that ${\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}]\leq Y_{t}^{x}$ $\mathbb{P}$-a.s.. Since $\tau\in{\mathcal{}T}_{t}$ is chosen some arbitrary, we have $V_{t}^{x}=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}]\leq Y_{t}^{x}.$
Step 2. We now claim that $Y_{t}^{x}\leq V_{t}^{x}$. Let $\tau_{t}^{*,x}\in{\mathcal{}T}_{t}$ be defined as in (2.5). Since $\int_{0}^{\tau_{t}^{*,x}}(Y_{s}^{x}-R(X_{s}^{x}))dK_{s}^{x}=0$ $\mathbb{P}$-a.s. (see Definition 2.6 (iii)) and $Y_{s}^{x}>R(X_{s}^{x})$ for all $s\in(0,\tau_{t}^{*,x})$ (by definition of $\tau_{t}^{*,x}$), it holds that
@@
By putting $\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds-\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}$ into the left-hand side of (6.4) and taking the conditional $g$-expectation ${\mathcal{}E}_{t}^{g}[\cdot]$, $\mathbb{P}$-a.s.,
${\mathcal{}E}_{t}^{g}\bigg[\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}r(X_{s}^{x})ds+e^{-\int_{t}^{\tau_{t}^{*,x}}\beta_{u}du}Y_{\tau_{t}^{*,x}}^{x}\bigg]$
(6.5) $\;=Y_{t}^{x}+{\mathcal{}E}_{t}^{g}\bigg[-\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}g(s,Z_{s}^{x})ds+\int_{t}^{\tau_{t}^{*,x}}e^{-\int_{t}^{s}\beta_{u}du}Z_{s}^{x}dB_{s}\bigg]=Y_{t}^{x},$
where the first equality follows from the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [coquet2002filtration, Lemma 2.1] and the second equality follows from the same arguments presented to show that the term $\operatorname{II}_{t}$ in (6.2) equals zero; see also Definition 2.1 (iii).
Last, since $Y_{\tau_{t}^{*,x}}^{x}=R(X_{\tau_{t}^{*,x}}^{x})$ $\mathbb{P}$-a.s. (by definition of $\tau_{t}^{*,x}$ in (2.5)), we have
$Y_{t}^{x}={\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau_{t}^{*,x}}]\leq V_{t}^{x}=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}_{t}^{g}[\operatorname{I}_{t}^{x;\tau}]\quad\mbox{$\mathbb{P}$-a.s.,}$
as claimed. Therefore, $\tau_{t}^{*,x}$ is optimal to (2.2). This completes the proof.
@@
Since $\alpha$ is uniformly bounded (noting that it has values only in $\{0,1\}$), one can deduce that the parameters of the BSDE (6.7) satisfies all the conditions given in [pardoux1990adapted, Section 3]. Hence, there exists a unique solution $(\widetilde{Y}_{t}^{x;N,\alpha},\widetilde{Z}_{t}^{x;N,\alpha})_{t\in[0,T]}\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ to the controlled BSDE (6.7).
@@
where we have used the property of ${\mathcal{}E}_{t}^{g}[\cdot]$ given in [coquet2002filtration, Lemma 2.1].
Moreover, by using the same arguments presented for the ${\mathcal{}E}^{g}$-martingale property in (6.2) (see the proof of Proposition 2.8) we can deduce that the conditional $g$-expectation appearing in the right-hand side of the above equals zero (i.e., the integrand therein is an ${\mathcal{}E}^{g}$-martingale). Hence the claim holds.
Step 2. It suffices to show that for every $t\in[0,T]$ $\mathbb{P}$-a.s., $Y_{t}^{x;N}=\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{Y}_{t}^{x;N,\alpha}.$ Indeed, it follows from Step 1 that for every $\alpha\in{\mathcal{}A}$ the parameters of the BSDE (6.7) satisfies the conditions given in [pardoux1990adapted, Section 3]. Furthermore, the parameters of the BSDE (2.7) also satisfies the conditions (see Remark 2.9 (i)).
@@
(6.8) $F^{x;N}_{t}(Y_{t}^{x;N},Z_{t}^{x;N})\geq\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{F}_{t}^{x;N,\alpha}(Y_{t}^{x;N},Z_{t}^{x;N}).$
@@
Therefore, an application of [el1997backward, Proposition 3.1] ensures the claim to hold.
Step 3. Lastly, it follows from [el1997backward, Corollary 3.3] that the process $\alpha^{*,x;N}\in{\mathcal{}A}$ is optimal for the problem given in Step 2., i.e., for all $t\in[0,T]$ $\operatorname*{ess\,sup}_{\alpha\in{\mathcal{}A}}\widetilde{Y}_{t}^{x;N,\alpha}=\widetilde{Y}_{t}^{x;N,\alpha^{*,x;N}}.$ This completes the proof.
@@ 6.2 Proof of results in Section 3
Let $N\in\mathbb{N}$ and $\lambda>0$ be given. We prove (i) by showing that the parameters of the BSDE (3.8) satisfy all the conditions given in [pardoux1990adapted, Section 3] to ensure its existence and uniqueness to hold.
@@
is strictly decreasing and $N$-Lipschitz continuous, we are able to see that for every $\omega\in\Omega$, $t\in[0,T]$, and $y,\hat{y}\in\mathbb{R}$
@@
From (6.9) and (6.11) and the definition of $\overline{F}^{x;N,\lambda}$ given in (3.7), it follows that the desired priori estimate of $\overline{F}^{x;N,\lambda}$ holds. Hence an application of [pardoux1990adapted, Theorem 3.1] ensures the existence and uniqueness of the solution of (3.8), as claimed.
@@
(6.12) $\overline{F}_{t}^{x;N,\lambda}(\omega,y,z)=F_{t}^{x}(\omega,y,z)+\max_{a\in[0,1]}\Big\{N(R(X_{t}^{x}(\omega))-y)a-\lambda{\mathcal{}H}(a)\Big\},$
@@
Therefore, an application of [el1997backward, Proposition 3.1] ensures the claim to hold.
Moreover, a direct application of [el1997backward, Corollary 3.3] ensures that $\pi^{*,x;N,\lambda}$ is optimal for $\overline{V}^{x;N,\lambda}$ given in (3.2). This completes the proof.
@@
Moreover, as the terminal conditions of (3.8) and (2.7) are coincide, it follows from the comparison principle of BSDEs (see, e.g., [el1997backward, Theorem 2.2]) that (3.10) holds.
@@
Since the parameters of the BSDEs (3.8) and (2.7) satisfy the conditions given in [el1997backward, Section 5] (with exponent $2$) for all $N\in\mathbb{N}$ and $\lambda>0$, we are able to apply [el1997backward, Proposition 5.1] to have the following a priori estimates: [^4] In [el1997backward, Section 5], the filtration (denoted by $({\mathcal{}F}_{t})$ therein) is set to be right-continuous and complete (and hence not necessarily the Brownian filtration, as in our case). Nevertheless, we can still apply the stability result given in [el1997backward, Proposition 5.1], since the martingales $M^{i}$, $i=1,2$, appearing therein are orthogonal to the Brownian motion. Consequently, the arguments remain valid when the general filtration is replaced with the Brownian one. for every $N\in\mathbb{N}$ and $\lambda>0$
@@
The following lemma establishes the monotone convergence of the hitting times of an increasing sequence of continuous processes, and it will serve as a key ingredient in the proof of Theorem 3.6.
Lemma 6.5.
Let $({L})_{t\in[0,T]}$, $(\mathcal{Y}_{t})_{t\in[0,T]}$, and $(\mathcal{Y}_{t}^{n})_{t\in[0,T]}$, $n\in\mathbb{N}$, be real-valued, $\mathbb{F}$-progressively measurable continuous processes satisfying that
  • (i)
    for every $n\in\mathbb{N}$ and $t\in[0,T]$, $\mathcal{Y}_{t}^{n}\leq\mathcal{Y}_{t}^{n+1}\leq\mathcal{Y}_{t}$ -$\mathbb{P}\text{-a.s.}$,
  • (ii)
    $\Delta^{n}:=\sup_{t\in[0,T]}(\mathcal{Y}_{t}-\mathcal{Y}_{t}^{n})\downarrow 0$ -$\mathbb{P}\text{-a.s.}$, as $n\uparrow\infty$.